Time value Personal Finance Problem You can deposit $10,000 into an account paying 13% annual interest either today or exactly 5 years from today. How much better off will you be at the end of 30 years if you decide to make the initial deposit today rather than 5 years from today? The future value at the end of 30 years if you deposit $10,000 at 13% today is (Round to the nearest dollar.) The future value at the end of 30 years if you deposit $10,000 at 13% 5 years from today is $ (Round to the nearest dollar.) If you invest the $10,000 now instead of waiting for 5 years to make the investment, you would be better off by $ (Round to the nearest dollar.)
Answer #1

1)

Future value if deposited today = Present value ( 1 + r)n

Future value if deposited today = 10,000 ( 1 + 0.13)30

Future value if deposited today = 10,000 * 39.115898

Future value if deposited today = $391,159

2)

Future value if deposited 5 years from today = Present value ( 1 + r)n

Future value if deposited 5 years from today = 10,000 ( 1 + 0.13)25

Future value if deposited 5 years from today = 10,000 * 21.230542

Future value if deposited 5 years from today = $212,305

3)

Better off by = 391,159 - 212,305

Better off by = $178,854